OpenStudy (anonymous):

How are you going to get the root of x^4 + 4?

OpenStudy (karatechopper):

help! me ishaan i dont get this

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

\[\sqrt{x^4+4}\]?

OpenStudy (anonymous):

by roots do you mean: x^4+4=0 ? You won't have any real roots.

OpenStudy (anonymous):

Oh sorry yea.=0. How bout imaginary roots?

OpenStudy (karatechopper):

imaginary? lol

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^4%2B4%3D0

OpenStudy (anonymous):

\[x = 4^{\frac{1}{4}}\] \[x = 2^\frac{1}{2}\]

OpenStudy (anonymous):

\[x =\sqrt2\]

OpenStudy (anonymous):

How did you do that? Haha

OpenStudy (anonymous):

shouldnt it be \[-4^{1/4}\]?

OpenStudy (anonymous):

Lol sorry fiddle is right I am trying to get the solution you will have to wait a little

OpenStudy (anonymous):

Okay! Thanks! :)

OpenStudy (anonymous):

\[x = (-1)^{\frac{1}{4}}*4^{\frac{1}{2}}\]

OpenStudy (karatechopper):

ishaan i have posted some songs on chat do u remember them?

OpenStudy (anonymous):

\[x = (-1)^{\frac{1}{4}}* \sqrt2\]

OpenStudy (anonymous):

Lol Okay karate let me check

OpenStudy (anonymous):

Now all we need to do is to find \((-1)^{\frac{1}{4}}\)

OpenStudy (phi):

\( -4^{\frac{1}{4}} \) in polar form is \( \sqrt{2}e^{\frac{i \pi} {4}}\) or equivalently \( \sqrt{2} ( cos(\pi/4) + i sin(\pi/4)) \) this simplifies to 1+i the other 3 roots are -1+i, -1-i, 1-i

OpenStudy (anonymous):

ah pellet i forgot the polar form

OpenStudy (anonymous):

nice work phi

OpenStudy (nikvist):

\[x^4+4=0\] \[x^4+4x^2+4-4x^2=0\] \[(x^2+2)^2-(2x)^2=0\] \[(x^2-2x+2)(x^2+2x+2)=0\] \[x^2-2x+2=0\quad\Rightarrow\quad x_{1,2}=\frac{2\pm\sqrt{4-8}}{2}=1\pm i\] or \[x^2+2x+2=0\quad\Rightarrow\quad x_{3,4}=\frac{-2\pm\sqrt{4-8}}{2}=-1\pm i\]

OpenStudy (anonymous):

nice tricks up there nikvist ^^^